Change the following equations in to augmented matrix

x-y+2z=1

3x-y+5z=-2

4x+2y+(x2-8)z=(x+2)


And determine values of x where:-

There is no solution

And where there is exactly one solution

And where there infinitely many solutions


1
Expert's answer
2021-05-24T12:29:30-0400

Augmented matrix 


"A=\begin{bmatrix}
 1 & -1 & 2 & 1 \\
 3 & -1 & 5 & -2 \\
4 & 2 & x^2-8 & x+2
\end{bmatrix}"



"R_2=R_2-3R_1"

"\begin{bmatrix}
 1 & -1 & 2 & 1 \\
 0 & 2 & -1 & -5 \\
4 & 2 & x^2-8 & x+2
\end{bmatrix}"



"R_3=R_3-4R_1"

"\begin{bmatrix}
 1 & -1 & 2 & 1 \\
 0 & 2 & -1 & -5 \\
0 & 6 & x^2-16 & x-2
\end{bmatrix}"



"R_2=R_2\/2"

"\begin{bmatrix}
 1 & -1 & 2 & 1 \\
 0 & 1 & -1/2 & -5/2 \\
0 &6 & x^2-16 & x-2
\end{bmatrix}"



"R_1=R_1+R_2"

"\begin{bmatrix}
 1 & 0 & 3/2 & -3/2 \\
 0 & 1 & -1/2 & -5/2 \\
0 & 6 & x^2-16 & x-2
\end{bmatrix}"



"R_3=R_3-6R_2"

"\begin{bmatrix}
 1 & 0 & 3/2 & -3/2 \\
 0 & 1 & -1/2 & -5/2 \\
0 & 0 & x^2-13 & x+13
\end{bmatrix}"

I) no solution


"\begin{matrix}
 x^2-13=0 \\
 x+13\not=0
\end{matrix}=>x=\pm\sqrt{13}"



Ii) exactly one solution


"x^2-13\not=0=>x\not=\pm\sqrt{13}"



Iii) infinitely many solutions


"\begin{matrix}
 x^2-13=0 \\
 x+13=0
\end{matrix}=>x\in \varnothing"